Nitrogen gas, N_{2} can be prepared by passing gaseous ammonia over solid copper(II) oxide, CuO, at high temperature. The other products of the reaction are solid copper, Cu, and water vapor.


2NH3 (g) + 3CuO (s) → → N₂ (g) + 3Cu (s) + 3H₂O(g)


In a certain experiment, a reaction mixture containing 100.0 g CuO and excess amount of ammonia, NH 3 , yields 7.46 g N_{2} What is the percent yield for this experiment?


1
Expert's answer
2022-02-06T11:20:30-0500

Reaction:

3CuO + 2NH3 = N2 + 3Cu + 3H2O

n(CuO) = "{\frac {m(CuO)} {64+16}}={\frac {90.4} {80}}=1.13" mol

n(NH3) = "{\frac {m(NH_3)} {14+3}}={\frac {18.1} {17}}=1.065" mol

For 1.065 mol of NH3 undergoing the reaction, "{\frac {3} {2}}" times larger amount of CuO is needed.

Needed amount of CuO is 1.6 mol but it's only 1.13 of it so CuO is a limiting reactant, we must count stoichiometry by amount of this reactant.

n(CuO) = 1.13 mol, n(N2 formed) = "{\frac {n(CuO)} {3}}={\frac {1.13} {3}}=0.3767" mol

m(N2) = 0.3767*28 = 10.55 grams